# Long Chain-of-Thought for Reasoning Large Language Models

### v4

@@ 8.5 Knowledge-Augmented Long CoT
The reasoning model significantly enhances reasoning capabilities, but it still lacks knowledge in specialized fields and timely new information [88, 165, 452, 650]. Thus, enriching reasoning with additional knowledge presents a key challenge for Long CoT [78, 71]. Current research focuses primarily on two approaches: (1) Retrieval-Augmented Generation: Retrieval-Augmented Generation (RAG) techniques enhance LLMs by integrating dynamic knowledge retrieval and document refinement [397, 781, 207, 303, 797, 1061, 1058, 566]. Research has combined RAG with reasoning modules to improve performance on complex tasks [698, 310, 451, 828, 83, 1020, 590]. O1 Embedder [884] optimizes multi-task retrieval and reasoning through synthetic data training. Furthermore, Stream of Search (SoS) [180], and CoRAG [757] boost search accuracy and addresses unresolved issues by incorporating more natural reflection and exploration in RAG. (2) Model Knowledge Injection: An alternative approach involves integrating additional knowledge during SFT or RL [473, 993, 119, 1090]. Specifically, HuatuoGPT-o1 [78] utilize the R1-like paradigm to train LLMs by model-judged reward RL, which significantly improves the medical knowledge during reasoning [551, 277]. Huang et al. [282] and Wang et al. [739] optimize for injecting medical knowledge in Long CoT scenarios by SFT, which also achieve great performance. Further, Jiang et al. [306] introduce MCTS to synthesize data, achieving superior performance. This model merges verifiable medical knowledge with reinforcement learning techniques to enhance performance in complex, medical task settings.
@@ 8.6 Safety and Stability for Long CoT
Despite the notable performance improvements brought about by Long CoT, Long CoT-augmented LLMs still encounter significant safety and stability challenges [1093, 1032, 492, 807]. These include issues such as the generation of unstable outputs, exemplified by the tendency to memorize in-domain math questions instead of engaging in actual reasoning [883], and the production of unsafe outputs, such as misinformation and offensive content [1081, 365, 1080, 487, 22, 42, 41, 150, 327, 1021]. Current research primarily addresses two key approaches: (1) Long CoT Attack Several studies show that Long CoT makes models more vulnerable to unexpected behavior [169, 139], hallucinations [239, 482] or unsafe outputs [341, 1102, 872, 103, 19, 502]. For instance, Arrieta et al. [23] identify that DeepSeek-R1 is prone to generating harmful content, including misinformation and offensive speech. Additionally, Kumar et al. [338] introduce the OverThink attack, which exploits false inference problems to induce overthinking in models, providing insights into potential defensive strategies. Further, Yao et al. [922] fool RLLMs chain of iterative chaos, for better jailbreaking. (2) Long CoT Safety Improvement Another major area of research focuses on enhancing safety [301, 1095, 470] and reliability [687, 610, 721, 140, 100, 628] through prompting [178] or training [553] techniques. Shen et al. [635] present Heima, which optimizes inference efficiency and robustness. Gallego [178] proposes dynamic security prompts during inference, while Cheng et al. [116] address hallucinations by guiding reasoning with a tree search algorithm. Zhao et al. [1050] introduce a self-reflection framework to identify biases, and Wang et al. [744] propose Safety Reasoning with Guidelines (SRG) to defend against out-of-distribution attacks. Finally, Parmar and Govindarajulu [561] combine reinforcement learning (RL) and supervised fine-tuning (SFT) in a hybrid training approach to reduce harmful outputs and enhance DeepSeek-R1’s safety.
@@ 9 Related Work
With the development of advanced RLLMs, such as OpenAI-o1 and DeepSeek-R1, recent research has focused on improving reasoning capabilities, especially on mathematical reasoning [765, 1054]. Patil [562] highlight the limitations of standard LLMs in addressing complex reasoning tasks, such as optimization and multi-step reasoning. In addition, Liang et al. [417] and Li [398] review strategies to scale search and inference time, including the use of algorithms like Monte Carlo Tree Search, to enhance LLM reasoning. Xu et al. [865] examine the role of reinforcement learning and "thought" sequences in reasoning improvement [340], while Hong et al. [242] demonstrate the impact of prompting techniques. Further, Liu et al. [450] and Mondorf and Plank [532] stress the importance of deeper analysis beyond surface-level accuracy, and He et al. [233] explore self-evolutionary processes as a means to advance LLM reasoning. Besta et al. [46] propose a modular framework integrating structure, strategy, and training methods as part of a comprehensive system design approach. Most recently, Li et al. [411] provide a systematic survey of System 2 thinking, focusing on the methods used to differentiate them from System 1 thinking.

### v5

@@ 8.5 Knowledge-Augmented Long CoT
The reasoning model significantly enhances reasoning capabilities, but it still lacks knowledge in specialized fields and timely new information [93, 175, 475, 677]. Thus, enriching reasoning with additional knowledge presents a key challenge for Long CoT [83, 75]. Current research focuses primarily on two approaches: (1) Retrieval-Augmented Generation: Retrieval-Augmented Generation (RAG) techniques enhance LLMs by integrating dynamic knowledge retrieval and document refinement [418, 811, 221, 322, 827, 1103, 1100, 592, 438]. Research has combined RAG with reasoning modules to improve performance on complex tasks [726, 329, 474, 861, 88, 1060, 616]. O1 Embedder [919] optimizes multi-task retrieval and reasoning through synthetic data training. Furthermore, Stream of Search (SoS) [193], and CoRAG [786] boost search accuracy and addresses unresolved issues by incorporating more natural reflection and exploration in RAG. (2) Model Knowledge Injection: An alternative approach involves integrating additional knowledge during SFT or RL [496, 1031, 124, 1132]. Specifically, HuatuoGPT-o1 [83] utilize the R1-like paradigm to train LLMs by model-judged reward RL, which significantly improves the medical knowledge during reasoning [577, 294, 769]. Huang et al. [300] and Wang et al. [766] optimize for injecting medical knowledge in Long CoT scenarios by SFT, which also achieve great performance. Further, Jiang et al. [325] introduce MCTS to synthesize data, achieving superior performance. This model merges verifiable medical knowledge with reinforcement learning techniques to enhance performance in complex, medical task settings.
@@ 8.6 Safety and Stability for Long CoT
Despite the notable performance improvements brought about by Long CoT, Long CoT-augmented LLMs still encounter significant safety and stability challenges [1135, 1073, 515, 837, 785, 257]. These include issues such as the generation of unstable outputs, exemplified by the tendency to memorize in-domain math questions instead of engaging in actual reasoning [918], and the production of unsafe outputs, such as misinformation and offensive content [1123, 384, 1122, 510, 23, 46, 45, 160, 346, 1061]. Current research primarily addresses two key approaches: (1) Long CoT Attack Several studies show that Long CoT makes models more vulnerable to unexpected behavior [181, 146], hallucinations [255, 505] or unsafe outputs [360, 1145, 906, 108, 20, 525]. For instance, Arrieta et al. [24] identify that DeepSeek-R1 is prone to generating harmful content, including misinformation and offensive speech. Additionally, Kumar et al. [357] introduce the OverThink attack, which exploits false inference problems to induce overthinking in models, providing insights into potential defensive strategies. Further, Yao et al. [958] fool RLLMs chain of iterative chaos, for better jailbreaking. (2) Long CoT Safety Improvement Another major area of research focuses on enhancing safety [320, 1138, 493] and reliability [715, 636, 748, 147, 105, 655] through prompting [191] or training [579] techniques. Shen et al. [662] present Heima, which optimizes inference efficiency and robustness. Gallego [191] proposes dynamic security prompts during inference, while Cheng et al. [121] address hallucinations by guiding reasoning with a tree search algorithm. Zhao et al. [1092] introduce a self-reflection framework to identify biases, and Wang et al. [772] propose Safety Reasoning with Guidelines (SRG) to defend against out-of-distribution attacks. Finally, Parmar and Govindarajulu [587] combine reinforcement learning (RL) and supervised fine-tuning (SFT) in a hybrid training approach to reduce harmful outputs and enhance DeepSeek-R1’s safety.
@@ 9 Related Work
With the development of advanced RLLMs, such as OpenAI-o1 and DeepSeek-R1, recent research has focused on improving reasoning capabilities, especially on mathematical reasoning [795, 1096, 33]. Patil [588] highlight the limitations of standard LLMs in addressing complex reasoning tasks, such as optimization and multi-step reasoning. In addition, Liang et al. [440] and Li [419] review strategies to scale search and inference time, including the use of algorithms like Monte Carlo Tree Search, to enhance LLM reasoning. Xu et al. [899] examine the role of reinforcement learning and "thought" sequences in reasoning improvement [359], while Hong et al. [259] demonstrate the impact of prompting techniques [546]. Further, Liu et al. [473] and Mondorf and Plank [557] stress the importance of deeper analysis beyond surface-level accuracy, and He et al. [248] explore self-evolutionary processes as a means to advance LLM reasoning. Besta et al. [50] propose a modular framework integrating structure, strategy, and training methods as part of a comprehensive system design approach. Most recently, Li et al. [432] provide a systematic survey of System 2 thinking, focusing on the methods used to differentiate them from System 1 thinking.